# RiemannLoRA: A Unified Riemannian Framework for Ambiguity-Free LoRA Optimization

### v1

@@ E.1 Ablation study on initialization
E.2 Math
In addition to the experiments on Commonsense reasoning, we fine-tuned Llama 3.2 1b on a $128$k subset ($2000$ steps with batch size equal to $64$) of the MetaMathQA dataset (see Section B.2). The accuracy is reported on the GSM8K evaluation set. Table 10 contains the accuracy of the trained via SGD without Heavy-Ball momentum model’s responses on the test dataset.
Table 10: The accuracy (in %) of fine-tuned Llama 3.2-1b using different SGD variants, tested on GSM8K and Commonsense Reasoning benchmarks. The <<R->> prefix stands for Riemannian, the postfix <<-LOI>> stands for locally optimal initialization. LoRA rank is set to $16$. In all experiments excpet for <<-LOI>> the $A$ factor in $AB^{\top}$ (for the first step) has orthonormal columns and $B$ is zero.
Setup GSM8K accuracy CR accuracy
LoRA $46.8$ $67.9$
LoRA-LOI $50.0$ $70.1$
DoRA $42.5$ $60.4$
RiemannLoRA $49.3$ $69.6$
RiemannLoRA-LOI $51.9$ $73.4$

### v2

@@ H.1 Ablation study on initialization
Appendix I Computational cost
To measure execution time, we used a virtual server instantiated on a compute node equipped with an Intel Xeon Gold 6240R CPU and Nvidia Tesla V100 GPU. The virtual machine was provisioned with 8 CPU cores, 16 GB of RAM and a single Tesla V100 GPU. Experiments were executed using the Hugging Face transformers library; timing corresponds to the execution time of the trainer.train() call. Measurements were conducted according to the following procedure. For each combination of method, LoRA adapter rank, and batch size, four timing runs were performed and the minimum execution time was selected. The measured quantity was the time spent on 16 optimizer steps. Time spent on initialization was excluded. For each combination of rank and batch size, execution times for each of the methods were measured sequentially. In the present work, rank refers to the rank of the LoRA adapter, and method denotes one of two optimization schemes: Adam (baseline) and Riemannion (proposed method).
Figure 4 shows the relative increase in per-step execution time of the proposed Riemannion method compared with Adam, computed as
Figure 4: Relative Time Cost $\left(\text{calculated as }\left(T_{\text{Riemannion}}-T_{\text{Adam}}\right)\big/{T_{\text{Adam}}}\right)$ of Riemannion vs. Adam during Llama 3-8B fine-tuning, as a function of LoRA rank and batch size.
Appendix J LLM usage
We used an LLM only for minor language polishing to improve readability and grammar; it was not involved in research ideation, methodology, analysis, or substantive writing, and all research ideas and arguments were developed entirely by the authors.